OpenStudy (anonymous):

poster w/ a perimeter of 100 inches. Length is 1.5 times the width. find the Area.

OpenStudy (anonymous):

2l + 2w = 100 l = 1.5w 3w + 2w = 100 w = 20 l = 1.5 *20 = 30 area = 20* 30 = 600 sq ins

OpenStudy (anonymous):

your the man thanks again

OpenStudy (anonymous):

no probs